8(4.50A plumber makes $4.50 per hour more than his apprentice. During an 8-hour day their combined earnings total $372. How much does each make per hour? (Hint: If you decide to use cents instead of dollars, then use 450 cents per hour and 37200 cents total earnings.)
Answer by checkley77(12844)

8(x+x+4.50)=372
8(2x+4.50)=372
16x+36=372
16x=372-36
16x=336
x=336/16
x=$21 per hour for the apprentice.
21+4.50=$25.50 for the plumber.
Proof:
8(21+25.50)=372
8*46.50=372
372=372